OpenStudy (janu16):

Rewrite in simplest radical form x 5 6 x 1 6. Show steps. I will draw if you don't understand what i wrote

OpenStudy (janu16):



OpenStudy (danjs):

hello again

OpenStudy (danjs):

this is the same two rules as before... what did you do first

OpenStudy (danjs):

to simplify, i wouldnt write any radicals

OpenStudy (janu16):

wait you do 5-1 so 4/6?

OpenStudy (danjs):

yes!

OpenStudy (janu16):

finally so you just put that in radical form?

OpenStudy (danjs):

\[x^{5/6} * x^{-1/6} = x^{5/6 + (-1/6)}\]

OpenStudy (danjs):

you can, i would just leave it as a fractional power,

OpenStudy (janu16):

its says rewrite in simplest radical form..

OpenStudy (danjs):

oh, right, yeah do that

jimthompson5910 (jim_thompson5910):

don't forget to reduce the fraction as much as possible

OpenStudy (janu16):

ohh we so would it be 1/3 or somethinf? @DanJS

OpenStudy (danjs):

4/6 = 2/3

OpenStudy (janu16):

oh ya 2/3 so is this how you show all the steps? hold on